A wave has an angular frequency of 110 rad/s and a wavelength of 1.80m. Calculate the
angular wave number and the speed of the wave
1
Expert's answer
2020-03-13T11:04:09-0400

"\omega =110 \ rad/s,\; \lambda =1.80 \ m"

"k, v \ - \ ?"

Solution:

Angular wave number: "k=\frac{2\pi}{\lambda}"

"k=\frac{2\pi}{1.80}=3.49 \ m^{-1}"


Speed of the wave: "v=\frac{\omega}{k}"

"v=\frac{110}{3.49}=31.51\ m/s"


Answer: "k=3.49 \ m^{-1},\; v=31.51 \ m/s."